OpenStudy (cookie_2046):

Solve the formula for the variable w. P = 2 l + 2w Question 12 options: w = (P/2) - l w = (P/2) + l w = P - l w = (P/2) - ( l / 2)

OpenStudy (dehelloo):

A strong suggestions would be https://www.youtube.com/watch?v=i3pfsCS7fWI

OpenStudy (cookie_2046):

dehelloo please stop

OpenStudy (dehelloo):

If you dont have yellow fever Id suggest just some australian footy like https://www.youtube.com/watch?v=SjHxXel0hZI

OpenStudy (dehelloo):

<3

OpenStudy (anonymous):

Sir this is the OS CoC enforcement stop immediately or we will send a fine directly to your house

OpenStudy (anonymous):

Heyo Cookie ^_^ need help?

OpenStudy (cookie_2046):

yes

OpenStudy (anonymous):

give me a sec

OpenStudy (cookie_2046):

k

OpenStudy (anonymous):

is that a one? after the 2

OpenStudy (dehelloo):

LOL I live in Australia. Seems legit that you wield the power to prosecute overseas.

OpenStudy (cookie_2046):

watch one are u talking about

OpenStudy (dehelloo):

<3 From an australian I'll leave you with one last video. https://www.youtube.com/watch?v=SjHxXel0hZI

OpenStudy (anonymous):

Oh Open Study has an HQ in every country ^_^

OpenStudy (anonymous):

the p = 2 1

OpenStudy (cookie_2046):

it looks like a variable

OpenStudy (anonymous):

It still doesn't make sense :/

OpenStudy (cookie_2046):

it doesnt look like it on here

OpenStudy (dehelloo):

OS is outsourced to India from an american company. Very small actually. Just saying mate. Only 1 location.

OpenStudy (anonymous):

looks very weird o-o

OpenStudy (cookie_2046):

looks like your not gonna help me...

OpenStudy (breanna1999):

First you have to divide 2 from the l on both sides

OpenStudy (breanna1999):

If I had to guess D would be the answer

OpenStudy (peridot):

@Cookie_2046 Don't fret, here's a variety of the equations and how it's simplified-- 2w = P - 2l w = (P-2l)/2